2x+3y=6
2x-5y=22
x=? y=?

Answer by oscargut(2103)

2x+3y=6
2x-5y=22
then
-2x-3y=-6
2x-5y=22
adding the equations
-8y=16
then y=-2
-2x-3(-2)= -6
-2x=-6-6
-2x=-12
x=6
Answer: x=6 and y=-2